OpenStudy (anonymous):

Simplify Cosx^2*Cscx^2-Cscx^2

OpenStudy (callisto):

Which one is your question, (i) or (ii)? \[(i) cosx^2\times cscx^2-cscx^2\]\[(ii) cos^2x\times csc^2x-csc^2x\]

OpenStudy (anonymous):

uhmm. actually it is theta... i put x to make it simpler to write. It is exactly Cos^2(theta)*csc^2(theta)-csc^2(theta)

OpenStudy (anonymous):

so i guess the 1st

OpenStudy (callisto):

\[cos^2\theta\times csc^2\theta-csc^2\theta\] This?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any ideas?

OpenStudy (callisto):

lol Then, it's the second :) \[cos^2\theta\times csc^2\theta-csc^2\theta\]\[ = csc^2\theta (cos^2\theta -1)\]\[=csc^2\theta [-(1-cos^2\theta )]\]\[=csc^2\theta (-sin^2\theta )\] One step left (or 2..), can you work it out?

OpenStudy (anonymous):

you got cos^2-1 from finding a GCF?

OpenStudy (callisto):

kinda.... Actually, by factorization

OpenStudy (anonymous):

Can you please show me it since its all ^2?

OpenStudy (callisto):

Hmm... actually, I've shown you... but I can explain little more... Consider \[cos^2\theta\times csc^2\theta-csc^2\theta\] \(csc^2\theta\)is the common factor of the 2 terms, got it so far?

OpenStudy (anonymous):

yse i think so

OpenStudy (anonymous):

so wouldnt it be csc^2(1-1)?

OpenStudy (anonymous):

if you just did the common factor

OpenStudy (callisto):

Nope :| Take OUT the common factor and group the rest of the terms...